solve each system using either methods substitution or elimination

2m+3n=4
-m+2n=5

Answer by stanbon(75887)

2m+3n=4
-m+2n=5
-------------
2m + 3n = 4
-2m + 4n = 10
----
7n = 14
---
n = 2
============
Solve for m:
2m + 3*2 = 4
2m + 6 = 4
2m = -2
m = -1
=============
Cheers,
Stan H.
=============